### v1

@@ 5.5 Higher dimensional models
In this final experiment we assess the accuracy of the barrier-constrained methods as the dimension of the state ($d_{x}$) and the observations ($d_{y}$) increase. Otherwise, the stochastic Lorenz 96 model parameters are $\sigma_{x}=\sigma_{y}=1$ and the continuous-time horizon is $T=10$, with observations collected every $\Delta_{o}=0.1$ continuous time units. For the barrier-based filters, we set $\mu=50$, $\kappa=100$ and $\varrho=4$.
To be specific, we compare the performance of 5 types of filters:
@@
  • a barrier-constrained standard PF (labeled barrier-SIR), and
  • a barrier-constrained EnKF (barrier-EnKF).
In Figure 3(a) we plot the NMSE attained by each filter versus the state dimension $d_{x}$ for a fixed number of particles (or ensemble size for the EnKFs) $N=750$. The observation dimension is set as $d_{y}=\lfloor 0.6d_{x}\rfloor$ and the pattern of observations is fixed for each simulation but chosen randomly across different simulations. The results displayed are an average of 10 independent runs.
We observe that the standard PFs break down quickly as the dimension increases and yield NMSEs over 1, which means that the power of the error is of the same order as the power of the signal $X_{n}$ itself. The performance of the standard EnKF degrades quickly when $d_{x}>N$, while both the barrier-constrained PF and the barrier-constrained EnKF yield stable NMSEs as $d_{x}$ increases. The NMSE attained by the barrier-EnKF filter is around half the error of the barrier-SIR filter.
Figure 3(b) shows the results of a similar experiment where the number of particles for the different algorithms is increased linearly with the model dimension, namely $N=d_{x}$. The simulation shows that the standard PFs still fail to track the stochastic Lorenz 96 model, while the barrier-based algorithms remain stable and improve their performance. The barrier-EnKF filter yields an NMSE which is consistently lower than the NMSE of the standard EnKF for $d_{x}\geq 40$, with a very similar computational cost.
@@
Figure 3: NMSE versus state-space dimension $d_{x}$ for SIR, EnKF, APF, barrier-SIR, and barrier-EnKF algorithms. (a) All filters run with a fixed sample size $N=750$. (b) Sample size is scaled with dimension as $N=d_{x}$.
@@ 6 Conclusions
We have investigated a class of constrained Bayesian filters for state space models where the source and target of the Markov transition kernels are compact subsets of the state space constructed from the observed data. In this setup, we have analyzed the stability and approximation errors of Bayesian filters. To be specific, we have proved that, under mild assumptions on the original (unconstrained) state-space model, the constrained filter is exponentially stable. Stability is an important feature because it enables the design of numerical filters which do not accumulate error over time. We have also studied the approximation errors of constrained filters and obtained error rates that depend on a parameter of the constraint. Our analysis also provides guidelines on the selection of the sequence of compact sets that form the constraint, namely, that they should be selected as superlevel sets of the likelihood functions in the state space model. The results of the approximation error analysis can also be leveraged to prove that particle filters run on the constrained state space model can be used to approximate the optimal (unconstrained) Bayesian filter under some regularity assumptions. Moreover, the selection of the compacts sets in the constraint as superlevel sets of the likelihood functions provides means to mitigate weight degeneracy in the particle filter.
In the second half of the paper we have turned our attention to the practical implementation of the proposed constrained filters in problems where data are collected over a discrete time grid but the signal of interest is a continuous-time stochastic Itô process characterised by a SDE. We have proposed a relatively simple approximation of the resulting constrained Markov kernels using barrier functions to modify the drift of the original SDE. As a result, we obtain a class of modified SDEs that can be easily simulated. Based on this approach, we have shown some numerical results that illustrate the stability and accuracy of the proposed constrained filters when tracking a (possibly high dimensional) stochastic Lorenz 96 system with partial and noisy observations.
@@ Acknowledgments
Let $f:{\mathcal{X}}\mapsto\mathbb{R}$ be a measurable test function, with $\|f\|_{\infty}<\infty$. The difference $\pi_{n}(f)-\pi_{n}^{N}(f)$ can be expanded as
$\pi_{n}^{N}(f)-\pi_{n}(f)$ $=$ $\sum_{i=0}^{n-1}\Phi_{n-i:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n-i:n}\left(\Phi_{n-i}\left(\pi_{n-i-1}^{N}\right)\right)(f)$
$+\Phi_{0:n}\left(\pi_{0}^{N}\right)(f)-\Phi_{0:n}\left(\pi_{0}\right)(f).$
and a straightforward application of Minkowski’s inequality yields
(31) $\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}$ $\leq$ $\sum_{i=0}^{n-1}\left\|\Phi_{n-i:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n-i:n}\left(\Phi_{n-i}\left(\pi_{n-i-1}^{N}\right)\right)(f)\right\|_{p}$
$+\left\|\Phi_{0:n}\left(\pi_{0}^{N}\right)(f)-\Phi_{0:n}\left(\pi_{0}\right)(f)\right\|_{p}.$
for any $p\geq 1$.
For $i=0$ in (31), we have
(32) $\left\|\Phi_{n:n}\left(\pi_{n-i}^{N}\right)(f)-\Phi_{n:n}\left(\Phi_{n}\left(\pi_{n-1}^{N}\right)\right)(f)\right\|_{p}=\|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\|_{p}$
and it is straightforward to show that
$\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)$ $=$ $\frac{\xi_{n}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}-\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\pm\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}$
$=$ $\frac{\xi_{n}^{N}(g_{n}f)-K_{n}\pi_{n-1}^{N}(g_{n}f)}{\xi_{n}^{N}(g_{n})}+\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\frac{K_{n}\pi_{n-1}^{N}(g_{n})-\xi_{n}^{N}(g_{n})}{\xi_{n}^{N}(g_{n})}.$
Since $\left|\frac{K_{n}\pi_{n-1}^{N}(g_{n}f)}{K_{n}\pi_{n-1}^{N}(g_{n})}\right|=|\Phi_{n}(\pi_{n-1}^{N})(f)|\leq\|f\|_{\infty}$, we readily see that
(33) $\left|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\right|\leq\frac{2\|f\|_{\infty}}{\xi_{n}^{N}(g_{n})}\left|\xi_{n}^{N}(g_{n})-K_{n}\pi_{n-1}^{N}(g_{n})\right|.$
From assumption (b) in the statement of Theorem 2.6 we obtain $\xi_{n}^{N}(g_{n})\geq\zeta>0$ and a straightforward application of the Marcinkiewicz-Zygmund inequality [^1] Note that $K_{n}\pi_{n-1}^{N}({\sf d}x_{n})=\frac{1}{N}\sum_{j=1}^{N}K_{n}(x_{n-1}^{j},{\sf d}x_{n})$ and $\xi_{n}^{N}({\sf d}x_{n})=\frac{1}{N}\sum_{j=1}^{N}\delta_{\bar{x}_{n}^{j}}({\sf d}x_{n})$, where $\bar{x}_{n}^{j}\sim K_{n}(x_{n-1}^{j},{\sf d}x_{n})$. If we let $\bar{\mathcal{F}}_{n-1}$ be the $\sigma$-algebra generated by $\{x_{0:n-1}^{1:N},\bar{x}_{1:n-1}^{1:N}\}$, it is clear that $\mathbb{E}[f(\bar{x}_{n}^{j})|\bar{\mathcal{F}}_{n-1}]=\int f(x_{n})K_{n}(x_{n-1}^{j},{\sf d}x_{n})$ for every j=1,,Nj=1,...,N. yields
(34) $\left\|\pi_{n}^{N}(f)-\Phi_{n}(\pi_{n-1}^{N})(f)\right\|_{p}\leq\frac{2c_{p}\|f\|_{\infty}\|g\|_{\infty}}{\zeta\sqrt{N}},$
where $c_{p}$ is a constant that depends only on $p$ (but is independent of $\xi_{n}^{N}$ and, in particular, of $N$) and we have used the bound $|g_{n}|\leq\|g\|_{\infty}$, also from assumption (b).
For $1\leq i\leq n-1$ in (31), let us first note that, from the mixing assumption (a) in Theorem 2.6 and Lemma 2.5, we arrive at
$\left\|\Phi_{n-i:n}(\pi_{n-i}^{N})-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))\right\|_{\rm tv}$ $\leq$
$\frac{(1-\gamma^{2})^{i-1}}{\gamma^{2}}\left\|\Phi_{n-i+1}(\pi_{n-i}^{N})-\Phi_{n-i+1}(\Phi_{n-i}(\pi_{n-i-1}^{N}))\right\|_{\rm tv}$ $=$
(35) $\frac{(1-\gamma^{2})^{i-1}}{\gamma^{2}}\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{n-i+1}(\pi_{n-i}^{N})(f)-\Phi_{n-i+1}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f).\right|$
However, for any $j\geq 2$, one can show that
(36) $\Phi_{j}(\pi_{j-1}^{N})(f)$ $=$ $\frac{\int f(x_{j})g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})\xi_{j-1}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})\xi_{j-1}^{N}({\sf d}x_{j-1})}$
$=$ $\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
where $K_{j}(g_{j}f)(x_{j-1})=\int f(x_{j})g_{j}(x_{j})K_{j}(x_{j-1},{\sf d}x_{j})$, and, similarly,
(37) $\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)$ $=$ $\frac{\int f(x_{j})g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}$
$=$ $\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}.$
Combining (36) and (37) yields
$\Phi_{j}(\pi_{j-1}^{N})(f)-\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)$ $=$
$\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}-\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}\pm\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$ $=$
$\frac{\xi_{j-1}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)-K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j}f)g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
(38) $+\rho_{j}(f)\frac{K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)-\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
where
$\rho_{j}({\sf d}x_{j})=\frac{g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}{\int g_{j}(x_{j})\int K_{j}(x_{j-1},{\sf d}x_{j})g_{j-1}(x_{j-1})K_{j-1}\pi_{j-2}^{N}({\sf d}x_{j-1})}$
is a probability measure. From (38) and the fact that $|\rho_{j}(f)|\leq\|f\|_{\infty}$ we readily find the inequality
(39) $\left|\Phi_{j}(\pi_{j-1}^{N})(f)-\Phi_{j}(\Phi_{j-1}(\pi_{j-2}^{N})(f)\right|\leq\frac{2\|f\|_{\infty}\left|\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)-K_{j-1}\pi_{j-2}^{N}\left(K_{j}(g_{j})g_{j-1}\right)\right|}{\xi_{j-1}^{N}\left(K_{j}(g_{j})g_{j-1}\right)}$
and, combining (35) and (39) with j=ni+1j=n-i+1 yields
$\sup_{\|f\|_{\infty}\leq 1}\left|\Phi_{n-i:n}(\pi_{n-i}^{N})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f)\right|$ $\leq$
(40) $\frac{2(1-\gamma^{2})^{i-1}}{\gamma^{2}}\frac{\left|\xi_{n-i}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)-K_{n-i}\pi_{n-i-1}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)\right|}{\xi_{n-i}^{N}\left(K_{n-i+1}(g_{n-i+1})g_{n-i}\right)}.$
Recalling assumption (b) of Theorem 2.6 ($\inf_{n}g_{n}\geq\zeta>0$) and the mixing assumption (a) ($\gamma u_{j}(x_{j})\leq k_{j}(x_{j-1},x_{j})\leq\frac{1}{\gamma}u_{j}(x_{j})$) we realise that
$|K_{n-i+1}(g_{n-i+1})g_{n-i}|\leq\frac{\|g\|_{\infty}}{\gamma}\int g_{n-i+1}(x_{n-i+1})u_{n-i+1}(x_{n-i+1}){\sf d}x_{n-i+1}$
and
$K_{n-i+1}(g_{n-i+1})g_{n-i}\geq\gamma\zeta\int g_{n-i+1}(x_{n-i+1})u_{n-i+1}(x_{n-i+1}){\sf d}x_{n-i+1}>0.$
Then, combining (40) with the Marcinkiewicz-Zygmund inequality, applied in the same way as in (34), we readily arrive at the $L^{p}$ bound
(41) $\left\|\Phi_{n-i:n}(\pi_{n-i}^{N})(f)-\Phi_{n-i:n}(\Phi_{n-i}(\pi_{n-i-1}^{N}))(f)\right\|_{p}\leq\frac{2c_{p}\|g\|_{\infty}(1-\gamma^{2})^{i-1}}{\gamma^{4}\zeta}\times\frac{1}{\sqrt{N}}.$
that holds for any measurable test function $f:{\mathcal{X}}\mapsto\mathbb{R}$ with $\|f\|_{\infty}\leq 1$.
The calculation for the last term on the right hand side of (31) is very similar. Lemma 2.5 yields
(42) $\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{0:n}(\pi_{0}^{N})(f)-\Phi_{0:n}(\pi_{0})(f)\right|\leq\frac{(1-\gamma^{2})^{n-1}}{\gamma^{2}}\sup_{\|f\|_{\infty}\leq 1}\frac{1}{2}\left|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right|$
and
$\left|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right|$ $\leq$
$\left|\frac{K_{1}\pi_{0}^{N}(g_{1}f)-K_{1}\pi_{0}(g_{1}f)}{K_{1}\pi_{0}(g_{1})}+\Phi_{1}(\pi_{0}^{N})(f)\frac{K_{1}\pi_{0}(g_{1})-K_{1}\pi_{0}^{N}(g_{1})}{K_{1}\pi_{0}(g_{1})}\right|$ $\leq$
(43) $\frac{2\|f\|_{\infty}}{K_{1}\pi_{0}(g_{1})}\left|K_{1}\pi_{0}^{N}(g_{1})-K_{1}\pi_{0}(g_{1})\right|$
and, using assumption (b) ($g_{1}>\zeta>0$) and the Marcinkiewicz-Zygmund inequality again, we arrive at
(44) $\left\|\Phi_{1}(\pi_{0}^{N})(f)-\Phi_{1}(\pi_{0})(f)\right\|_{p}\leq\frac{2\|f\|_{\infty}\|g\|_{\infty}c_{p}}{\zeta\sqrt{N}},$
where the constant $c_{p}<\infty$ depends only on $p\geq 1$. Hence, for any measurable and real test function $f$ with $\|f\|\leq 1$, the combination of (42), (43) and (44) yields
(45) $\left\|\Phi_{0:n}(\pi_{0}^{N})(f)-\Phi_{0:n}(\pi_{0})(f)\right\|_{p}\leq\frac{2c_{p}\|g\|_{\infty}(1-\gamma^{2})^{n-1}}{\gamma^{2}\zeta\sqrt{N}}.$
Finally, taking together (31), (34), (41) and (45) yields
$\sup_{\|f\|_{\infty}\leq 1}\|\pi_{n}^{N}(f)-\pi_{n}(f)\|_{p}\leq\frac{2c_{p}\|g\|_{\infty}}{\gamma^{6}\zeta}\times\frac{1}{\sqrt{N}}.\quad\Box$
Appendix B Proof of Theorem 3.3
@@
(53) $\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)$ $=$ $\frac{\xi_{n}(g_{n}f)}{\xi_{n}(g_{n})}-\frac{\hat{\xi}_{n}^{l}(g_{n}f)}{\hat{\xi}_{n}^{l}(g_{n})}\pm\frac{\hat{\xi}_{n}^{l}(g_{n}f)}{\xi_{n}(g_{n})}$
$=$ $\frac{\xi_{n}(g_{n}f)-\hat{\xi}_{n}^{l}(g_{n}f)}{\xi_{n}(g_{n})}+\hat{\pi}_{n}^{l}(f)\frac{\hat{\xi}_{n}^{l}(g_{n})-\xi_{n}(g_{n})}{\xi_{n}(g_{n})},$
@@
Appendix C Proof of Theorem 3.5
Let $\Phi_{n}$ be the PU operator associated to model $\mathcal{S}$} and let $\Phi_{i:j}=\Phi_{j}\circ\cdots\circ\Phi_{i+1}$, with $\Phi_{i:i}(\pi)=\pi$ for any measure $\pi\in{\mathcal{P}}({\mathcal{X}})$. With this notation, the approximation error can be written as
@@
Appendix D Proof of Theorem 3.6
@@
Appendix E Proof of Theorem 3.7

### v2

@@ 5.5 Higher dimensional models
In this final experiment we assess the accuracy of the barrier-constrained methods as the dimension of the state ($d_{x}$) and the observations ($d_{y}$) increase. Otherwise, the stochastic Lorenz 96 model parameters are $\sigma_{x}=\sigma_{y}=1$ and the continuous-time horizon is $T=10$, with observations collected every $\Delta_{o}=0.1$ continuous time units. For the barrier-based filters, we set $\mu=50$, $\kappa=100$ and $\varrho=4$. The schemes are rather robust to the choice of $\kappa$, however $\mu$ needs to be increased for moderate and large $d_{x}$.
To be specific, we compare the performance of 7 types of filters:
@@
  • a barrier-constrained standard PF (Barrier-SIR)
  • a barrier-constrained EnKF (Barrier-EnKF)
  • a guided PF (Guided) implemented as in [Pieper-Sethmacher25b], and
  • a barrier-constrained (Barrier-Guided) that incorporates a barrier term to the guided PF of [Pieper-Sethmacher25b].
For the guided PF we employ Algorithm 1 in [Pieper-Sethmacher25b], with the auxiliary drift chosen as $a(t)=0$. Under this choice, the simplified auxiliary dynamics reduces to a linear Ornstein–Uhlenbeck process without deterministic forcing. Although alternative choices of $a(t)$, such as frozen or filtered-mean linearizations, are possible and may further improve performance, we restrict attention here to the zero-drift auxiliary model in order to provide a clear baseline assessment of the method. The barrier-guided PF is constructed by combining the guidance mechanism with the barrier correction. Concretely, the proposal drift is obtained by augmenting the original Lorenz–96 drift with the guidance term and subsequently adding the barrier-induced correction term. The weight update remains that of Algorithm 1, so that the barrier modification acts only at the level of the proposal dynamics. This construction enables us to examine the effects of guidance and barrier stabilization in high-dimensional settings.
In Figure 3(a) we plot the NMSE attained by each filter versus the state dimension $d_{x}$ for a fixed number of particles (or ensemble size for the EnKFs) $N=750$. The observation dimension is $d_{y}=\lfloor 0.6d_{x}\rfloor$ and the observation matrix $H$ is fixed for each simulation but chosen randomly across different simulations. The results displayed are an average of 128 independent simulation runs.
We observe that the standard PFs break down quickly as the dimension increases and yields NMSEs over 1, which means that the power of the error is of the same order as the power of the signal $X_{n}$ itself.The guided PF improves upon the standard PF but still deteriorates as the dimension $d_{x}$ grows, with NMSEs remaining significantly above those of the barrier-based methods. The performance of the standard EnKF degrades quickly when $d_{x}>N$, while both the barrier-constrained PF and the barrier-constrained EnKF yield stable NMSEs as $d_{x}$ increases. The barrier-guided PF further improves the guided method and achieves consistently lower NMSEs than the simpler guided PF. The NMSE attained by the barrier-EnKF filter is around half the error of the barrier-SIR filter.
Figure 3(b) shows the results of a similar experiment where the number of particles for the different algorithms is increased linearly with the model dimension, namely $N=\lfloor 0.75d_{x}\rfloor$. The simulation shows that the standard PFs still fail to track the stochastic Lorenz 96 model, while the barrier-based algorithms remain stable and improve their performance. The guided PF improves upon the standard PF under this scaling but still exhibits a gradual increase in NMSE as the dimension grows. The barrier-guided filter further improves the performance of the guided PF and maintains stable NMSE levels across dimensions. The barrier-EnKF filter yields an NMSE which is consistently lower than the NMSE of the standard EnKF for $d_{x}\geq 40$, with a very similar computational cost.
Finally in Table 1 we illustrate that the three barrier-based methods we consider in Figure 3(a) are applicable to larger dimensions. Here computations are made with a fixed sample size of $N=750$.
@@
Figure 3: NMSE versus state-space dimension $d_{x}$ for SIR, EnKF, APF, barrier-SIR, and barrier-EnKF algorithms. (a) All filters run with a fixed sample size $N=750$. (b) Sample size is scaled with dimension as $N=\lfloor 0.75d_{x}\rfloor$.
$d_{x}$ 2500 5000 7500
Barrier-SIR 0.2560 0.2705 0.1604
Barrier-ENKF 0.1214 0.1485 0.2757
Barrier-Guided 0.2770 0.2762 0.2797
Table 1: NMSE for the three Barrier based methods for $d_{x}=2500,5000,7500$. The results are averaged over four independent runs.
@@ 6 Conclusions
We have investigated a class of constrained Bayesian filters for state space models where the source and target of the Markov transition kernels are compact subsets of the state space constructed from the observed data. In this setup, we have analyzed the stability and approximation errors of Bayesian filters. To be specific, we have proved that, under mild assumptions on the original (unconstrained) state-space model, the constrained filter is exponentially stable. Stability is an important feature because it enables the design of numerical filters which do not accumulate error over time. We have also studied the approximation errors of constrained filters and obtained error rates that depend on a parameter of the constraint. Our analysis also provides guidelines on the selection of the sequence of compact sets that form the constraint, namely, that they should be selected as superlevel sets of the likelihood functions in the state space model.
In the second half of the paper we have turned our attention to the practical implementation of the proposed constrained filters in problems where data are collected over a discrete time grid but the signal of interest is a continuous-time stochastic Itô process characterised by a SDE. We have proposed a relatively simple approximation of the resulting constrained Markov kernels using barrier functions to modify the drift of the original SDE. As a result, we have obtained a class of modified SDEs that can be easily simulated and built into sampling-based numerical filters, including particle filters and ensemble Kalman filters. Based on this approach, we have shown some numerical results that illustrate the stability and accuracy of the proposed constrained filters when tracking a (possibly high dimensional) stochastic Lorenz 96 system with partial and noisy observations.
@@ Acknowledgments
(33) $\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)=\frac{\xi_{n}(g_{n}f)}{\xi_{n}(g_{n})}-\frac{\hat{\xi}_{n}^{l}(g_{n}f)}{\hat{\xi}_{n}^{l}(g_{n})}\pm\frac{\hat{\xi}_{n}^{l}(g_{n}f)}{\xi_{n}(g_{n})}=\frac{\xi_{n}(g_{n}f)-\hat{\xi}_{n}^{l}(g_{n}f)}{\xi_{n}(g_{n})}+\hat{\pi}_{n}^{l}(f)\frac{\hat{\xi}_{n}^{l}(g_{n})-\xi_{n}(g_{n})}{\xi_{n}(g_{n})},$
@@
Appendix B Proof of Theorem 3.5
Let $\Phi_{n}$ be the PU operator associated to model $\mathcal{S}$} and let $\Phi_{i:j}=\Phi_{j}\circ\cdots\circ\Phi_{i+1}$, with $\Phi_{i:i}(\pi)=\pi$ for any measure $\pi\in{\mathcal{P}}({\mathcal{X}})$. The approximation error can be written as
@@
Appendix C Proof of Theorem 3.6
@@
Appendix D Proof of Theorem 3.7